Integration by parts

1. Integral of 5 sin^-1 (x) dx

2. Integral of 2 lnx dx
1
Expert's answer
2019-12-11T10:35:29-0500

Calculus

We need to find the Integration by parts.


Solution:


We know the formula of Integration by Parts


"\int f(x) \space g(x) \space dx = f(x) \space \int g(x) dx - \int f'(x) (\int g(x) dx) \space dx"


1).


"\int 5 sin^{-1} x dx" = 5 "\int sin ^{-1} x \times 1 \space dx"


Here, f(x) = "sin^{-1} x" and g(x) = 1 ( By Using ILATE)


"\int 5 sin^{-1} x dx" = 5 [ "sin^{-1} x \int 1 dx - \int d(sin^{-1} x) (\int 1 dx) dx" ]


"= 5 [ x sin^{-1} x - \int \frac {1}{\sqrt {1-x^2}}\times x dx ]"


"= 5 [x sin^{-1} x - \frac {1}{2} \int \frac {2x}{\sqrt {1-x^2}} dx]"


"= 5 [ x sin^{-1} x + \frac {1}{2} \int t^ {\frac {-1}{2}} dt] + c"

Here "t = 1 - x^2 \\
dt = - 2x dx"

"\int 5 sin^{-1} x dx" =

"5 [ x sin^{-1} x + \frac {1}{2} \frac {t^{\frac {-1}{2} + 1}}{\frac {-1}{2}+1}] + c"

"= 5 [ x sin^{-1} x + t^{ \frac {1}{2}}] + c = 5 [ x sin^{-1} x + \sqrt {1-x^2}] +c"

2).


"\int 2 ln x = 2 \int ln x \times 1 \space dx"


Using ILATE , let f(x) = ln x and g(x) = 1



"\int 2 ln x = 2 [ln x \int 1 dx- \int f' (x) (\int g(x) dx )dx]"

"= 2 [ x \space ln x - \int \frac {1}{x} \times x dx ] + c"

"= 2 [ x \space ln x - x ] + c"